OpenStudy (anonymous):

Water tank rate problem? How to represent represent h as a function of t if dV/dt = 50 = pi(h)^2(k)^2dh/dt? (Water is flowing at a rate of 50m^3/min, the cone shaped tank is 10m in height, and the diameter is 40m.)

OpenStudy (anonymous):

Please, someone help me with this! I it is the last problem on my review and I really want to get the concept down.

OpenStudy (anonymous):

\[V=1/3*\pi*r^2*h\] \[r:h=2:1\] \[r=2h\] \[V=1/3*\pi*4h^2*h=4\pi h^3/3\] I'm sorry if I am too late.

OpenStudy (anonymous):

\[dV/dt=4\pi h^2dh/dt\] \[50=\pi h^2dh/dt \] \[50/\pi h^2=dh/dt\]